OpenStudy (gitface):

If a rock is thrown upwards from level ground, and its max height is equal to its horizontal range, R, what angle was the rock thrown?

OpenStudy (gitface):

my diagram is;  My current equations are; \[v _{y} = 0, \Delta _{y}=\frac{ 1 }{ 2 }R\] So, using,\[\Delta _{y}=v _{o}t+\frac{ 1 }{ 2 }at ^{2}\] I get; \[y _{\max}=v _{oy}\frac{ 1 }{ 2 }t+\frac{ 1 }{ 2 }a \frac{ 1 }{ 2 }t ^{2}\] Becasue \[\Delta _{y} = \frac{ 1 }{ 2 }R\]